OpenStudy (anonymous):

The sum of the measures of the angles of a triangle is 180 . The largest angle measures 10° less than the sum of the 2 others, the measure of the middle size angle is the average of the other 2 find the measures of the 3 angles

OpenStudy (anonymous):

let the angles be x,y,z and x>y>z x+y+z=180 ...(1) x=y+z-10 y+z=x+10 x+x+10=180 2x=170 x=85 y+z=180-85=95 ...(2) also \[y=\frac{ x+z }{ 2 },2y=85+z,z=2y-85 ...(3)\] put in (2) and solve and complete.